The hypotenuse of a right triangle is 6 more than twice the shortest side. If the third side is 2 m less than the hypotenuse, find the sides of the triangle.

Grade:12th Pass

1 Answers

Ritvik Gautam
12 years ago

Let the shortest side be x

Then , hypotenuse = (x+6)

and the third side =(x+6)-2 = (x+4)

By applying pythagoras theorm,

(x+6)2 = (x+4)2 + x2

x2 - 4x - 20=0

Hence, we can find the value of x and hence all the three sides.

 

 

Please Approve!